OpenStudy (anonymous):

how do you find the radius with 2 pi r

sam (.sam.):

do you know the area of the circle?

OpenStudy (anonymous):

the area of z=124.44 the area of sector XZY is 28 square

OpenStudy (anonymous):



OpenStudy (zepp):

This whole circle's area would become Area Z + Area XYZ: the area of z=124.44 the area of sector XZY is 28 square 124.44+28=152.44

OpenStudy (zepp):

Apply your formula, \(2 \pi r=152.44\) and solve for r.

OpenStudy (anonymous):

how would I solve for r (lost)

OpenStudy (zepp):

\(2* \pi* r = 152.44\) \(r = 152.44 \div 2\pi\)

OpenStudy (anonymous):

24.26 ?

OpenStudy (zepp):

Yep.

OpenStudy (anonymous):

thank you